OpenStudy (anonymous):

Please help unable to solve! (5^3)^2 x 5^0 -------------- What I get stuck on is the list of answers as the answer I get does not match up 1. 0 2. 5^5 3.5^6 4.5^30

OpenStudy (anonymous):

so 5^3=125

OpenStudy (anonymous):

125 ^ 2 = 15625

OpenStudy (anonymous):

5^0 = 1

OpenStudy (anonymous):

15625 x 1 = =)

OpenStudy (anonymous):

When I tried to complete the question, it turned to be 5300 But when it' does not appear in the list that's what i'm stuck on

OpenStudy (anonymous):

Oh I see your reply i'll look at it..

OpenStudy (anonymous):

15625? I got that when I used a calculator before yes.. But It does not show up as one of the options?

OpenStudy (anonymous):

Unless i'm missing it entirely..

OpenStudy (anonymous):

5^6?